OpenStudy (kikuo):

4x^2-28x+49=0 I need to factor this, but with with quadratic formula. Can you show me the steps and the answer? Will medal.

OpenStudy (isaidavila):

a=4 b=-28 c=49

OpenStudy (kikuo):

@isaidavila can you show how you got your answer?

OpenStudy (mathstudent55):

Use the quadratic formula. First, write the quadratic formula. Then rewrite it substituting 4 for a, -28 for b, and 49 for c. What do you get?

OpenStudy (isaidavila):

Just have to replace those values I gave you in the formula

OpenStudy (mathmale):

You are given 4x^2-28x+49=0 and need to pick out the coefficient of each of the 3 terms on the left. That's where the {4, -28, 49} comes from.

OpenStudy (mathmale):

Please type out (or draw) the quadratic formula, so I can be certain that you know it.

OpenStudy (kikuo):

@mathstudent55 ax^2+bx+c=0 4x^2-28x+49=0

OpenStudy (kikuo):

@isaidavila Is this correct?

OpenStudy (mathmale):

Yes. Notice how the first coefficient is a = 4, the second b = -28 and the third c = 49. tO FIND THE roots of this quadratic equation, we substitute these values of a, b and c in to the quadratic formula:\[x=roots=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (mathmale):

Note that -b in this equation becomes -(-28) = + 28.